# LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs

### v2

@@ Abstract.
Large Language Models (LLMs) have demonstrated impressive performance across a wide range of NLP tasks, including summarization. LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – remains largely unexplored. LLMs have a limited context window size, restricting the amount of data that can be processed at once. We tackle this challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries from large collections of user-generated text using LLMs. LaMSUM integrates summarization with different voting methods to achieve robust summaries. Extensive evaluation using four popular LLMs (Llama 3, Mixtral, Gemini, GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods. Overall, this work represents one of the first attempts to achieve extractive summarization by leveraging the power of LLMs, and is likely to spark further interest within the research community.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanyisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Brevity is the soul of wit.
— William Shakespeare, Hamlet, Act 2, Scene 2
Social media platforms like Facebook, X (formerly Twitter), and Reddit offer a medium for individuals to express their opinions and views on various subjects, leading to a diverse array of perspectives shared through social debates, critiques, and reviews  (Zhang et al., 2024b; Chhikara et al., 2023; Dash et al., 2019a; Bražinskas et al., 2020; Zhang et al., 2016). With a vast amount of data being generated by users online, there is a growing need for summarization algorithms providing a precise and concise summary, eliminating the need for users to sift through numerous posts or reviews. Based on the resulting summary, summarization algorithms can be categorized as ‘extractive’ and ‘abstractive’. In extractive summarization, the aim is to select a subset representative of the original text (Xu et al., 2020; Liu and Lapata, 2019; Zhong et al., 2020; Zhang et al., 2022, 2023a). In contrast, abstractive summarization aims to generate natural language summaries that capture the essence of the original text, often paraphrasing the content rather than directly extracting it (Lewis et al., 2020; Zhang et al., 2020).
Extractive summarization is a crucial task with a rich body of literature, widely applied in summarizing legal cases, news articles, lectures, clinical notes, and social media content (Bhattacharya et al., 2021; Kanwal and Rizzo, 2022; Mukherjee et al., 2020; Jia et al., 2020). For example, a Google search for a topic or hashtag displays a few tweets alongside the usual lists of websites and news articles. The selection of tweets displayed in search results is similar to an extractive summarization task, where only a subset of tweets is selected (Griffin, 2015). Additionally, the Library of Congress only stores a selection of tweets as part of its archive to optimize storage space (Congress, 2017). E-commerce platforms like Amazon display only a subset of reviews in the condensed view, such that the chosen reviews are representative of all the reviews submitted by the users. Such cases are instances of extractive summarization, where the target is to select the most relevant subset that can effectively summarize a topic or a product. When summarizing user generated data, it is crucial to quote the user’s exact words, making extractive summarization particularly valuable in such cases. In the process of summarization, each post or review is treated as a basic unit – often referred to as a textual unit.
@@ 1. Introduction
In recent years, Large Language Models (LLMs) have exhibited high performance across various tasks, including summarization (Ouyang et al., 2022; Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by the human evaluators (Pu et al., 2023; Liu et al., 2024). However, using LLMs for extractive summarization has two significant limitations: (i) as generative models, LLMs naturally tend to perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences; and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text. Given the increasing reliance on LLMs for summarization, it is worth investigating the utility of LLMs for extractive summarization of large volumes of user-generated text.
To address these limitations, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al., 2016). Voting algorithms are traditionally employed in decision making processes to aggregate preferences and opinions of a large population; the prime example being the electoral system, where voters’ preferences for candidates are gathered to elect a winner that best represents the voters’ preferences. Voting systems can be categorized into single-winner and multi-winner types, depending on the number of winners. In a single-winner system, only one candidate is elected, whereas a multi-winner system allows for the selection of multiple candidates.
Extractive summarization can be viewed as a multi-winner election, where the input units act as candidates, and the units chosen for the summary are considered as the winners. A multi-winner voting system includes approval-based committee (ABC) voting, where voters approve a subset of candidates without ranking, and ranked choice voting, where voters rank candidates by preference. In LaMSUM, we utilize two approval based voting algorithms – Plurality Voting (Mudambi et al., 1996) and Proportional Voting (Lackner et al., 2023), and one ranked based voting algorithm – Borda Count (Emerson, 2013). Our judicial application of voting algorithms with a multi-level summarization framework ensures that LaMSUM outperforms the state-of-the-art fine-tuned summarization models. In summary, in this work, we make the following contributions:
  • We propose a novel framework LaMSUM which can effectively summarize large (having ¿30K tokens) collection of user generated content.
@@
  • LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine outputs to generate robust summaries.
  • Analyse whether an ensemble model with multiple LLMs in LaMSUM can lead to improved outcomes across different voting methods.
To our knowledge, this is the first work to implement extractive summarization of large user-generated texts using LLMs by combining summarization with voting algorithms. We believe this work can spawn further research in this direction.
2. Background and Related Work
@@
Figure 1. Current LLMs, by default, produce abstractive summaries. Llama3-70b-8192, despite specifically prompted for extractive summarization, generates abstractive summaries. This behavior underscores the need for a targeted approach to enable LLMs to effectively generate extractive summaries.
Text Summarization Algorithms
With the growing volume of online data, the demand for algorithms that automatically shorten and summarize texts is increasing. Automatic summarization can be approached in two ways: extractive and abstractive. In extractive summarization, a subset of input collection is selected for the summary based on their perceived quality and significance, aiming to represent a larger dataset with a concise sample. Abstractive summarization, on the other hand, generates natural language summaries that capture the most critical information from the original text. Both approaches seek to provide readers with a concise overview of the textual content. Over the years, many text summarization algorithms have been proposed in the literature; the reader can refer to (Gupta and Lehal, 2010; El-Kassas et al., 2021) for detailed surveys.
Extractive Summarization of User Generated Content
A large body of prior research has focused on summarizing long documents (Allahyari et al., 2017; Dong, 2018; Cajueiro et al., 2023). In fact, traditionally, summarization tasks have focused on summarizing a single document, such as a news article or a business report. In recent years, summarization has been increasingly applied on different types of user generated text (e.g., tweets, Facebook or Reddit posts) (Mackie et al., 2014; Inouye and Kalita, 2011; Jiang et al., 2021), where the task is to summarize short, independent posts written by many users. Several extractive summarization algorithms tailored specifically for user generated content have also been proposed (Sharifi et al., 2010; Xu et al., 2013; Dash et al., 2019b; Mukherjee et al., 2022; Rudra et al., 2018; Jones et al., 2023).
@@
LLMs are now being extensively used for summarization (Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang et al., 2023b; Tang et al., 2023a; Yang et al., 2023; Bražinskas et al., 2020; Laskar et al., 2023), but their primary focus remains on short documents that can fit in the LLM context window. Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al., 2024a; Wu et al., 2024; Goyal et al., 2023; Zhang et al., 2023c; Liu et al., 2024). Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al., 2024; Tam et al., 2023; Luo et al., 2023; Laban et al., 2023).
Figure 2. LaMSUM: Multi-level framework for extractive summarization of large user-generated text. Input set $\mathcal{T}$ (level 0) is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks each of size s. From each chunk a summary is produced of size $q$ (refer Figure 3), $q$ length summaries from $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks are merged to form the input for the next level i.e., level 1. Iteratively the same procedure is repeated till we obtain a summary of size $k$. We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk (Section 3.2).
Extractive Summarization through LLMs: The Current State By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise. To illustrate this, we present a small example in Figure 1. An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 100 sentences, it fails to do so and instead generated an abstractive summary. Prior to our current work, only two studies attempted to perform similar tasks. Zhang et al. (2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al. (2024) attempted abstractive summarization for book-length documents. However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias in LLMs.
@@
3. Generating Extractive Summaries through LLMs
@@ 3.1. Task Formulation
3.2. Multi-Level Summarization
@@
LLMs have a limited context window, making it impossible to input large text collections all at once. While recent models like GPT-4 support context windows of up to 128k tokens, they still cannot accommodate book-length inputs within a single window. Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al., 2024). Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2)  [^1] Code is available at https://anonymous.4open.science/r/LaMSUM/ .
@@
The set $\mathcal{T}$, which contains the original textual units, is provided as input at level 0 and is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ number of chunks of size $s$. From each chunk of size $s$, we generate a summary (discussed in Section 3.3) of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks.  [^2] Note that a chunk of size $s$ refers to a chunk containing $s$ textual units. Likewise, a summary of size $q$ indicates a summary of $q$ textual units. $|\mathcal{T}|$ denotes the number of textual units present in $\mathcal{T}$. We then merge all these $q$ length summaries obtained from level 0 to form an input for the next level i.e., level 1. We repeatedly perform this process until we obtain the final summary of length $k$. Note that the last chunk maybe less than $q$ in size, in such case we move all the textual units of the respective chunk to the next level (refer Algorithm 1).
@@
An alternate strategy would be to divide the input $\mathcal{T}$ into $\frac{|\mathcal{T}|}{s}$ chunks each of size $s$ and from each chunk select $\frac{k\cdot s}{|\mathcal{T}|}$ sentences to be included in the summary. However, this approach assumes a uniform distribution of potential candidates across chunks that can be included in the final summary. In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate (discussed in Section 4.4). In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively.
@@
It is important to note that we are dealing with user-generated content, such as tweets, which lack contextual connections. Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, tweets are generally standalone and contextually independent. Thus, our approach of independently deriving summaries from each chunk works well in our setup, as each textual unit operates independently of the others and there are no long-range dependencies.
@@
Figure 3. Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias. $m$ different chunk variations are obtained through shuffling, which are subsequently summarized using LLMs. $m$ summaries are then aggregated by voting algorithms to get the final summary.
3.3. Summarizing a Chunk
@@
Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al., 2016).
3.3.1. Tackling Positional Bias
@@
Prior research  (Brown and Shokri, 2023; Zhang et al., 2023b; Jung et al., 2019; Wu et al., 2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary. To address this issue and generate a robust summary, we create $m$ different variations by shuffling the textual units within the input chunk. This ensures that each unit has the opportunity to appear in different positions within the input text (refer Figure 3).
3.3.2. Zero-shot prompting
@@
For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM. We employ the following two prompts to obtain the summaries –
a
@@
Select the most suitable units that summarize the input text.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that selects best ¡summary_length¿ sentences (subset) which summarizes the input. Think step by step and follow the instructions. ¡sentences¿”
b
Generate a ranked list in descending order of preference.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that outputs the sentences in the decreasing order of their relevance to be included in the summary. Think step by step and follow the instructions. ¡sentences¿ Remember that output should contain all the sentences in the decreasing order of their preference.”
@@
              $ei=|\mathcal{T}|$triangleright ending index is equal to length of $\mathcal{T}$
@@
         $width=ei-si$triangleright number of textual units in a chunk
@@
         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with random state $i$
         $R=\textsc{LLM}(V,q)$triangleright obtain summary from the LLM
@@
     return $\textsc{VOTING}(X,q)$triangleright voting to obtain the final summary
@@
  My allergies be acting up &amp; Claritin be letting them &gt;,&lt; My allergies be acting up & Claritin be letting them.
I spent the extra money to have fast relief, but instead I am SUFFERING #gimmemymoneyback I spent the extra money to have fast relief, but instead I am SUFFER
@@
Table 1. Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
3.3.3. Output Calibration
@@
LLMs may alter certain words from the input text while generating extractive summaries, as shown in Table 1. Thus, we perform additional checks to ensure that the textual units selected in the summary are indeed a subset of $\mathcal{T}$. If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos, 1998). LLMs may also hallucinate, generating new sentences rather than selecting units from the input. In such instances, the edit distance between the generated unit $x$ and all the original textual units would be high. We discard these sentences from the output.
3.4. Reimagining Summarization as an Election
@@
As mentioned earlier, for a given chunk, we obtain $m$ summaries – one for each variation. We imagine the process of creating the final summary from these $m$ summaries to be a multi-winner election, where the textual units in $m$ summaries correspond to ballots (candidates) and the role of the voting algorithm is to pick $q$ winners. We employ three different voting methods, namely Plurality Voting (Mudambi et al., 1996), Proportional Approval Voting (PAV) (Lackner et al., 2023) and Ranked Choice Voting (Emerson, 2013) to determine the final summary. Due to the varying input requirements of different voting methods, changing both the prompting approach and the output generated by the LLM becomes imperative.

### v4

# LaMSUM: Amplifying Voices Against Harassment through LLM Guided Extractive Summarization of User Incident Reports

@@ Abstract
Citizen reporting platforms like Safe City in India help the public and authorities stay informed about sexual harassment incidents. However, the high volume of data shared on these platforms makes reviewing each individual case challenging. Therefore, a summarization algorithm capable of processing and understanding various Indian code-mixed languages is essential. In recent years, Large Language Models (LLMs) have shown exceptional performance in NLP tasks, including summarization. LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – through LLMs remains largely unexplored. Moreover, LLMs have a limited context window size, restricting the amount of data that can be processed at once. We tackle these challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries for large collections of Safe City posts using LLMs. LaMSUM integrates summarization with different voting methods to achieve robust summaries. Extensive evaluation using three popular LLMs (Llama, Mistral and GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods for Safe City posts. Overall, this work represents one of the first attempts to achieve extractive summarization through LLMs, and is likely to support stakeholders by offering a comprehensive overview and enabling them to develop effective policies to minimize incidents of unwarranted harassment.
Warning: This paper contains content that may be disturbing or upsetting.
Introduction
  Category Post
  Robbery This incident took place in the evening.Two bikers came on a bike and snatched Rs.17000 from an old lady at gun point.
Stalking I was stalked by a guy who followed me for days and also he sent me letters on my doorstep saying he was madly in love with me and he is obsessed with my body.
Sexual Invites I was walking on footpath to a place nearby to meet my friends. A man was driving his car on parallel road and was continuously passing vulgar comments. I ignored but 15 min later he stopped his car and said “chalri h ky, paise h mere pas” (I have money, you want to come with me?). I screamed and asked for help from people around me.
Mas*****tion in public One day I reached my school prior to the school timings mistakenly. One of the drivers called me and my friend and started mas*****ting in front of us.
Ogling I was going to my coaching by driving my scooty when suddenly some boys came on a bike. They started making cheap comments and teasing me. It was late and the area was secluded. I got scared as they started revolving their bike around me. I started driving in the direction of a crowded place that is when they left.
Showing Po**ography A man in a car parked outside was watching po** when I saw him. He turned his device towards me and did offensive hand gestures inviting me.
Sexual Assault When I was 7 years old, the shopkeeper removed my clothes and started touching me everywhere. He also tried to do it to another girl and failed.
Domestic       Violence My husband always doubts on my character and doesn’t allow me to go outside alone, uses very vulgar language and beats me.
Table 1: Examples of harassment cases shared on the Safe City platform. Proactive action by authorities and citizens can help prevent numerous such incidents. Providing stakeholders with a concise overview of incidents occurring in a specific area is crucial and this can be effectively achieved by utilizing summarization algorithms.
In recent decades, the widespread availability of the internet has provided seamless access to online platforms to millions of people. Governments worldwide are increasingly utilizing these platforms to gather information directly from citizens – referred as Citizen Reporting (Kopackova and Libalova 2019). By leveraging tools such as mobile applications, web-based portals, and social media integrations, citizen reporting platforms establish a direct and efficient communication link between individuals and the relevant authorities, enabling faster issue resolution and facilitating active public participation in community improvement. Beyond immediate problem solving, real-time data gathered through these platforms contribute valuable information for urban planning and proactive measures, paving the way for more efficient and adaptive communities. Citizen reporting typically addresses topics such as community issues, environmental challenges, crime prevention, public health, and disaster response (Shin et al. 2024).
A notable citizen reporting platform is Safe City, developed by the nonprofit ‘Red Dot Foundation’ [^1] https://webapp.safecity.in; https://reddotfoundation.in, where people post incidents of sexual harassment, violence and assault. Table 1 showcases some example incidents shared by users on the platform. Although the platform is accessible worldwide via website and mobile app, the majority of its users are based in India. Despite making significant strides in recent decades in areas like Science & Technology, Defense and Agriculture, India continues to grapple with the critical issue of women safety. As per the ‘2023 Women Peace and Security Index’, India was positioned at 128th among 177 evaluated countries (Buchholz 2024). Violence against women remains an enduring and urgent concern, taking various forms, including domestic abuse, sexual harassment, rape, dowry-related violence, honor killings, and human trafficking. One such recent incident that deeply affected the nation was the rape and murder of a 31-year-old doctor while she was on duty in August 2024 (Times 2024). Several other publicized incidents include the 2020 Hathras case, where a young Dalit woman was brutally attacked and raped, ultimately resulting in her death (Dublish 2020) and the 2012 Nirbhaya incident involved the horrific gang rape and fatal assault of a 22-year-old woman in a private bus (Today 2020).
While such horrific incidents cannot be entirely avoided through reporting alone, platforms like Safe City can play a crucial role in preventing certain cases of sexual assault. By enabling users to analyze reported incidents, assess the safety of specific locations, and make informed decisions when traveling to potential hotspots, these platforms contribute to enhanced personal safety and awareness. Similarly, the local authorities can also benefit from these platforms to assess emerging cases, identify the underlying factors and determine proactive measures for effective resolution. However, the challenge for the authorities is to navigate the high volume of information in such platforms. Manually reviewing all posts is often impractical, necessitating a need for a summarization algorithm that can identify and select posts that are diverse as well as representative of the original data. Additionally, platforms like Safe City often feature a curated selection of posts on their homepage to showcase their core purpose, mission, and services. This deliberate selection also acts as a form of summarization.
Summarization algorithms are of two types: ‘extractive’ and ‘abstractive’. In extractive summarization, the algorithm selects a subset representative of the original text (Xu et al. 2020; Zhong et al. 2020; Zhang, Liu, and Zhang 2022, 2023a). In contrast, abstractive summarization algorithms generate summaries that capture the essence of the original text, often paraphrasing the content (Pu, Gao, and Wan 2023). For platforms like Safe City, extractive summarization is more suitable, as the goal is not to paraphrase the posts but to select a few that accurately capture a snapshot of the original content. When summarizing such sensitive posts, preserving the user’s exact words is essential, making extractive summarization particularly valuable in maintaining authenticity and context.
Several extractive summarization algorithms for user generated content have been proposed in the literature, primarily for text written in English (Bhattacharya et al. 2021; Kanwal and Rizzo 2022; Mukherjee et al. 2020; Jia et al. 2020). However, the Safe City platform receives posts in multiple languages from across India [^2] India is a linguistically rich and diverse country, home to 22 scheduled languages and many more spoken dialects., including numerous code-mixed entries where multiple languages are blended, such as Hinglish (a mix of Hindi and English). Such multilinguality limits the applicability of existing algorithms for extractive summarization of posts submitted to Safe City.
@@ Introduction
In recent years, Large Language Models (LLMs) have demonstrated very good performance across various tasks in multilingual and code-mixed settings (Ouyang et al. 2022; Brown et al. 2020; Tang et al. 2023a; Jin et al. 2024b). Plus, summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by human evaluators over other baseline algorithms (Pu, Gao, and Wan 2023; Liu et al. 2024). These prior results motivated us to investigate the utility of LLMs for extractive summarization of large volumes of user generated posts. However, we encountered two significant limitations which hinder the immediate application of LLMs for extractive summarization: (i) as generative models, LLMs perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences (as shown in Figure 1) (Worledge, Hashimoto, and Guestrin 2024); and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text (Jin et al. 2024a).
To overcome these limitations, in this paper, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al. 2016). Our judicial application of voting algorithms with a multi-level summarization framework ensures that LaMSUM outperforms the state-of-the-art fine-tuned summarization models. In summary, in this work, we make the following contributions:
@@
  • We propose a novel framework LaMSUM which can produce extractive summaries from large (having $>$30K tokens) collection of user generated content. LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine LLM outputs to generate robust summaries.
  • Extensive experiments demonstrates that LaMSUM outperforms the state-of-the-art extractive summarization algorithms.
  • We apply LaMSUM to user posts on the Safe City platform and develop a companion website that offers citizens and authorities a quick, comprehensive summary of harassment incidents occurring in localities across India.
To our knowledge, this is the first work to implement extractive summarization of large collection of user-generated texts using LLMs by combining summarization with voting algorithms. At the same time, we demonstrate the effectiveness of such algorithms to facilitate data-driven decision-making promoting safer communities by providing actionable insights into reported incidents. We believe this work can spawn further research in this direction, and to enable it, we are making the dataset, including the human-annotated gold standard summaries, available upon request. Code is available at https://anonymous.4open.science/r/LaMSUM/
Background and Related Work
@@
Figure 1: Current LLMs, by default, produce abstractive summaries. Llama-3.3-70b, despite specifically prompted for extractive summarization, generates abstractive summaries. This behavior underscores the need for a targeted approach to enable LLMs to effectively generate extractive summaries.
AI Solution through Citizen Reporting
Web and social media platforms receive posts on sensitive issues such as online harassment, hate speech, abusive behavior, violence etc. Abuse experienced by users lead to mental stress often forcing them to leave the platform (Sambasivan et al. 2019; Thomas et al. 2022; Kim et al. 2024). Several AI-powered solutions have been designed to address solutions to these critical issues. Machine learning based classifiers and language models are utilized to detect the cases of sexual abuse, hate speech, offensive language, human trafficking and harassment cases (Sawhney et al. 2021; Hassan et al. 2020; Davidson et al. 2017; Upadhayay, Lodhia, and Behzadan 2021; Stoop et al. 2019; Ghosh Chowdhury et al. 2019). Modelling the cyberbullying behavior using social network and language-based features can improve the classifier performance (Ziems, Vigfusson, and Morstatter 2020; Olteanu et al. 2018). Development of a mobile computing based reporting tool empowers individuals with intellectual and developmental disabilities (I/DD) to self report abuse and share the incident with the intended group (Venkatasubramanian et al. 2021; Sultana et al. 2021). With the rise of public figures encouraging women to speak up, the number of non-anonymous self-reported assault stories have increased (ElSherief, Belding, and Nguyen 2017). Counterspeech is proven to be a viable alternative to blocking or suspending problematic messages or accounts, as it better aligns with the principles of free speech (Mathew et al. 2019). Conversational agents (CAs) have attracted significant interest as potential counselors due to their features, such as anonymity, which can help address many challenges associated with human-human interaction (Park and Lee 2021).
@@
LLMs are now being extensively used for summarization (Brown et al. 2020; Tang et al. 2023a; Jin et al. 2024b). Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang, Liu, and Zhang 2023b; Tang et al. 2023b; Yang et al. 2023; Bražinskas, Lapata, and Titov 2020; Laskar et al. 2023), but their primary focus remains on short documents that can fit in the LLM context window. Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al. 2024; Wu et al. 2024; Goyal, Li, and Durrett 2023; Zhang, Liu, and Zhang 2023c; Liu et al. 2024). Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al. 2024; Tam et al. 2023; Luo, Xie, and Ananiadou 2023; Laban et al. 2023).
Figure 2: LaMSUM: Multi-level framework for extractive summarization of large user-generated text. Input set $\mathcal{T}$ (level 0) is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks each of size s. From each chunk a summary is produced of size $q$ (refer Figure 3), $q$ length summaries from $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks are merged to form the input for the next level i.e., level 1. Iteratively the same procedure is repeated till we obtain a summary of size $k$. We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk.
Extractive Summarization through LLMs: The Current State
By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise. To illustrate this, we present a small example in Figure 1. An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 50 sentences, it fails to do so and instead generated an abstractive summary. Prior to our current work, only two studies attempted to perform similar tasks. Zhang, Liu, and Zhang (2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al. (2024) attempted abstractive summarization for book-length documents. However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias.
@@
LaMSUM: Generating Extractive Summaries through LLMs
@@
Multi-Level Summarization
@@
LLMs have a limited context window, making it impossible to input large text collections all at once. While recent models like GPT-4 support context windows of up to 128k tokens, they still cannot accommodate book-length inputs within a single window. Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al. 2024). Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2). .
@@
The set $\mathcal{T}$, which contains the original textual units, is provided as input at level 0 and is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ number of chunks of size $s$. From each chunk of size $s$, we generate a summary of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks. [^3] Note that a chunk of size $s$ refers to a chunk containing $s$ textual units. Likewise, a summary of size $q$ indicates a summary of $q$ textual units. $|\mathcal{T}|$ denotes the number of textual units present in $\mathcal{T}$. We then merge all these $q$ length summaries obtained from level 0 to form an input for the next level i.e., level 1. We repeatedly perform this process until we obtain the final summary of length $k$. Note that the last chunk may be less than $q$ in size, in such case we move all the textual units of the respective chunk to the next level (refer Algorithm 1).
@@
An alternate strategy would be to divide the input $\mathcal{T}$ into $\frac{|\mathcal{T}|}{s}$ chunks each of size $s$ and from each chunk select $\frac{k\cdot s}{|\mathcal{T}|}$ sentences to be included in the summary. However, this approach assumes a uniform distribution of potential candidates across chunks that can be included in the final summary. In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate. In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively, as we keep $q=k$.
@@
It is important to note that we are dealing with user-generated content, such as posts, which lack contextual connections. Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, posts are generally standalone and contextually independent. Thus, our approach of independently deriving summaries from each chunk works well in our setup, as each textual unit operates independently of the others and there are no long-range dependencies.
@@
Figure 3: Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias. $m$ different chunk variations are obtained through shuffling, which are subsequently summarized using LLMs. $m$ summaries are then aggregated by voting algorithms to get the final summary.
Summarizing a Chunk
@@
Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al. 2016).
Tackling Positional Bias
@@
Prior research  (Brown and Shokri 2023; Zhang, Liu, and Zhang 2023b; Jung et al. 2019; Wu et al. 2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary. To address this issue and generate a robust summary, we create $m$ different variations by shuffling the textual units within the input chunk. This ensures that each unit has the opportunity to appear in different positions within the input text (refer Figure 3).
Zero-shot prompting
@@
For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM. We employ the following two prompt strategies to obtain the summaries (detailed in Appendix Table 7) –
1
@@
Select most suitable units that summarize the input text.
2 Generate a ranked list in descending order of preference.
@@
              $ei=|\mathcal{T}|$triangleright ei is equal to length of $\mathcal{T}$
@@
         $width=ei-si$triangleright no. of textual units in a chunk
@@
         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with state $i$
         $R=\textsc{LLM}(V,q)$triangleright obtain LLM summary
@@
     return $\textsc{VOTING}(X,q)$triangleright voting for final summary
@@
  In a train some people were staring me continuously. It was very uncomfortable Some people were staring me continuously in a train.
We were going to metro station, a biker started following us. When we shouted, he rode away A biker followed us and rode away when we shouted.
Some boy do dirty comments on me and my religion. Boys made dirty comments about my religion.
@@
Table 2: Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
Output Calibration
@@
LLMs may alter certain words from the input text while generating extractive summaries, as shown in Table 2. Thus, we perform additional checks to ensure that the textual units selected in the summary are indeed a subset of $\mathcal{T}$. If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos 1998).
LLMs may also hallucinate, generating new sentences rather than selecting units from the input. In such cases, the edit distance between the generated unit $x$ and the original textual units tend to be high. To address this, we adopt an alternative approach – we extract key elements such as nouns, verbs, and adjectives from the newly generated sentence $x$. Then search for the exact same set of keywords with the textual units in $\mathcal{T}$ and identify the post with the highest number of matching keywords as the closest match to $x$ (refer Algorithm 4 in Appendix).
Reimagining Summarization as an Election
@@
As mentioned earlier, for a given chunk, we obtain $m$ summaries – one for each variation. We imagine the process of creating the final summary from these $m$ summaries to be a multi-winner election, where the textual units in $m$ summaries correspond to ballots (candidates) and the role of the voting algorithm is to pick $q$ winners. We employ three different voting methods, namely Plurality Voting (Mudambi, Navarra, and Nicosia 1996), Proportional Approval Voting (PAV) (Lackner, Regner, and Krenn 2023) and Ranked Choice Voting (Emerson 2013) to determine the final summary. Due to the varying input requirements of different voting methods, changing both the prompting approach and the output generated by the LLM becomes imperative.